OpenStudy (anonymous):

Rebekah purchased a sweater that was reduced by twenty-five percent (or one-fourth) of its original price. She also used a store coupon for five dollars. When she made the purchase, she paid the cashier nineteen dollars. What was the original price of the sweater? twenty-four dollars twenty-eight dollars thirty-two dollars thirty-six dollars

OpenStudy (anonymous):

$32 32/4=8 32-8=24 24-5=19